A car moving at 15 m/s crashes into a tree and stops in 0.3 s. Calculate the force the seat belt exerts on a passenger in the car to bring him to a halt. The mass of the passenger is 88 kg. (Enter the magnitude.)
_N
Answer #1

here,

the initial speed of car , u = 15 m/s

final speed of car , v = 0 m/s

time taken , t = 0.3 s

let the acceleration be a

using first equation of motion

v = u + a * t

0 = 15 + a * 0.3

solving for a

a = - 50 m/s^2

mass of passenger , m = 88 kg

the magnitude of force the seat belt exerts on a passenger in the car to bring him to a halt , F = m * a

F = |88 * (-50)| N = 4400 N